OpenStudy (anonymous):

If a plane can travel 460 miles per hour with the wind and 400 miles per hour against the wind find the speed of the plane without a wind and the speed of the wind.

OpenStudy (anonymous):

v + wind = 460 v - wind = 400

OpenStudy (anonymous):

Okay to get the answer I would have to subtract the 460 first or divide it?

OpenStudy (anonymous):

To get the answer you have to solve for one variable in terms of the other, and plug into the second equation, or add the two equations to cancel out one variable.... So for the lower equation rearrange it... \[v - wind = 400\] can be rewritten as \[v = 400 + wind\] now plug this expression of v into the top equation!

OpenStudy (anonymous):

Still a bit confusing

OpenStudy (anonymous):

Do you understand how I have rewritten it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, so you want to plug this into the other equation where v is... \[v + wind = 460\] \[(400 + wind) + wind = 460\] \[2wind + 400 = 460\] \[2wind = 60\] \[wind = \frac{ 60 }{ 2 } = 30\]

OpenStudy (anonymous):

Okay got it.

OpenStudy (anonymous):

Solve the following two equaitons for s and w, the speed of the plane and the wind respectively: \[\{s+w=460,s-w=400\} \]\[\{s=430,w=30\} \]